Problem 13-4A Calculation of financial statement ratios LO P3 Selected year-end financial statements of Cabot Corporation follow. (All sales were on credit, selected balance sheet amounts at December 31, 2016, were inventory, $53,900; total assets, $259,400; common stock, $83,000; and retained earnings, $57,668.) CABOT CORPORATION Income Statement For Year Ended December 31, 2017 Sales Cost of goods sold Gross profit Operating expenses Interest expense Income before taxes Income taxes Net income $ 456,600 297,950 158,650 98,600 4,400 55,650 22,418 $ 33,232 CABOT CORPORATION Balance Sheet December 31, 2017 Liabilities and Equity Assets Cash Short-term investments Accounts receivable, net Notes receivable (trade) Merchandise inventory $ 22,000 Accounts payable $ 17,500 3,200 2,900 8,600 Accrued wages payable 29,000 Income taxes payable 7,000 40,150 Long-term note payable, secured by 63,400 mortgage on plant assets Prepaid expenses Plant assets, net Total assets 2,850 Common stock 151,300 Retained earnings 83,000 90,900 $ 260,900 $260,900 Total liabilities and equity * These are short-term notes receivable arising from customer (trade) sales Required: Compute the following:(1) current ratio, (2) acid-test ratio, (3) days sales uncollected, (4) inventory turnover, (5) days sales in inventory (6) debt-to-equity ratio, (7) times interest earned, (8) profit margin ratio, (9) total asset turnover, (10) return on total assets, and (11) return on common stockholders equity. (Do not round intermediate calculations.)Req 1 and 2 Req 3 Req 4 Req 5 Req 6 Req 7 Req 8 Req 9 Req 10 Req 11 Compute the current ratio and acid-test ratio urrent Ratio Choose Numerator: Choose Denominator:Current Ratio ccounts receivable net Accounts receivable Current ratio 2017: id-Test Rati Choose Numerator: minator: Acid-Test Ratio Acid-Test Ratio 2017: 0 to 1 Req 1 and 2 Req 3 >Req 1 and 2 Req 3 Req 4 Req 5 Req 6Req 7 Req 8 Req 9 Req 10 Req 11 Compute the days sales uncollected ays Sales Uncollected Choose Numerator: | 1 | Choose Denominator: | x Days | - | Days Sales Uncollected Days sales uncollected 2017: 0 days Req 1 and 2 Req 4 XReq 1 and 2 Req 3 Req 4 Req 5 Req 6 Req 7 Req 8 Req 9 Req 10 Req 11 Compute the inventory turnover Inventory Turnover Choose Numerator:Choose Denominator: I Inventory Turnover Inventory turnover 2017: 0 times Req 3 Req5 〉4Req 5Rq 7 Req 8 Req 1 and 2 Req 3 Req 9 Req 10 Req 11 Compute the days sales in inventory. Da Choose Numerator: Choose Denominator:x DaysDays Sales in Inventory Days sales in inventory 2017: 0 days K Req 4 Req 6 >Req 1 and 2 Req 3 Req 4Req 5 Req 6 Req 7 Req 8Req 9 Req 10 Req 11 Compute the debt-to-equity ratio. Debt-to-Equity Ratio Choose Numerator: Choose Denominator:Debt-to-Equity Ratio Debt-to-equity ratio 0 to 1 2017: Req 5 Req 7 >Req 1 and 2 Req 3 Req 4 Req 5 Req 6 Rq 7 Req 8 Req 9 Req 10Req 11 Compute the times interest earned. Times Interest Earned Choose Numerator: Choose Denominator: Times Interest Earned = - Times interest earned 2017: 〈 Req6 Req 8 〉Req 1 and 2 Req 3 Req 4Req 5 Req 6 Req 7 Req 8 Req 9Req 10 Req 11 Compute the profit margin ratio. Profit Margin Ratio Choose Numerator: Choose Denominator:Profit margin ratio Profit margin ratio 2017: 01%) 〈 Req7 Req 9 〉Req 1 and 2 Req 3 Req 4Req 5Req 6 Req 7 Req 8Req 9 Req 10 Req 11 Compute the total asset turnover Choose Numerator: I Choose Denominator:Total Asset Turnover Total asset turnover 2017: 0 times 〈 Req 8 Req 10 〉Req 1 and 2 Req 3 Req 4 Req 5 Req 6 Req 7 Req 8 Req 9Req 10 Req 11 Compute the return on total assets (10) Choose Numerator: Choose Denominator: Return on Total Assets Return on total assets 2017: 01% 〈 Req 9 Req 11Req 1 and 2 Req 3 Req 4 Req 5 Rq Rq 7 Req 8 Req 9 Req 10Req 1 Compute the return on common stockholders equity eturn on Co mmon Stockholders Equity Choose Numerator: Choose Denominator Return On Common Stockholders Equity Return on common stockholders equity 2017: Req 10 Req 11
Answer #1

1) current ratio = current assets / current liabilities

= 109600 / 23600

= 4.64

Note:- current assets = cash + short term investment + accounts receivable + notes receivable + merchandise inventory + perpaid expense

current liabilities = Accounts payable + accrued wage payable + income tax payable

2) current ratio = [current assets - merchandise inventory - perpaid expense ] / current liabilities

= 66600 / 23600

= 2.82

4) inventory turnover ratio = cost of goods sold / average inventory

= 297950 / [(53900 + 40150)/2]

= 297950 / 47025

= 6.34

5) days sales in inventory = number of days / inventory turnover ratio

= 365 / 6.34

= 57.57

6) debt to equity ratio = total liabilities / total equity

= 87000 / 173900

= 0.50

7) times interest earned = Earning before interest and tax / interest expense

=[income before tax + interest] / interest expense

= [55650 + 4400] / 4400

= 60050 / 4400

= 13.65

8) profit margin ratio = net income / sales

= 33232 / 456600

= 7.28%

9) total assets turnover = sales / average total assets

=456600 / [(259400 + 260900 )/2]

=456600 / 260150

= 1.76

10) return on total assets = net income / average total assets

= 33232 / [(259400 + 260900 )/2]

= 33232 / 260150

=12.77%

11) return on common stockholder's equity = net income / average common stockholder's equity

= 33232 / [({83000+57668}+{83000+90900})/2]

= 33232 / [(140668+173900)/2]

= 33232 / 157284

= 21.13%